OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x – 9y = –36 A. y = -4/9x - 4 B. 9y = 4x + 36 C. 9y = -4x - 36 D. y = 4-9x + 4

OpenStudy (anonymous):

standard form of the slope intercept form is y=mx+c where m=slope c=y intercept

OpenStudy (anonymous):

y=mx+b so, y= -4/9 +4

OpenStudy (raden_zaikaria):